# Domain Generalization for Medical Image Analysis: A Survey

### v2

@@ 3.2 Feature-level Generalization
Refer to Table III for a summary of feature-level methods. In the following paragraphs, we explore feature-level domain generalization techniques.
3.2.1 Feature Alignment
@@ 3.2.1.1   Feature Normalization
3.2.1.2   Dissimilarity-based Alignment
@@ 3.2.2 Disentanglement Methods
3.2.2.1   Implicit Feature Disentanglement
@@
Contrastive Disentanglement aims to make representations of similar instances more alike (low contrast) and those of different instances more dissimilar (high contrast). A typical contrastive learning loss function [77] is defined as:
@@
where $\operatorname{sim}(\cdot,\cdot)$ is the cosine similarity, $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are positive pairs, $k\in a(i)$ are the indexes of selected negative samples, and $\mathbf{x}_{i}$ and $\mathbf{x}_{k}$ are negative pairs. In the context of domain generalization, the positive pairs can be defined as samples from the same domain, while negative pairs are selected from other domains.
@@
3.2.2.2   Explicit Feature Disentanglement
@@
3.2.3 Other Representation Learning Methods
3.2.3.1   Feature Augmentation
@@
3.2.3.2   Kernel-based Learning
Kernel-based methods are a classic and effective approach within feature-level domain generalization. They operate by mapping the original input features into a higher dimensional space, often termed the feature space, where the data might be linearly separable or more structured, enabling better generalization across domains. There are various kernel-based methods for feature-level domain generalization, including Support Vector Machine (SVM) variants, Maximum Mean Discrepancy (MMD), and Transfer Component Analysis (TCA). Kernel trick enables these methods to operate in high-dimensional spaces without explicitly calculating the coordinates of the data in that space, but by simply computing the dot products between the images of all data pairs in the feature space. This makes the calculations more tractable and efficient. These kernel-based methods can benefit medical image analysis as they can handle high-dimensional data and discover complex patterns. They also offer an excellent way to incorporate domain knowledge, such as spatial relationships in images, by defining appropriate kernels.
@@ 3.2.3.2   Kernel-based Learning
3.3 Model-level Generalization
Model-level generalization serves as a principal technique among the plethora of strategies used to mitigate the challenges posed by domain shift, especially in the context of the medical imaging domain. In this sense, various strategies for promoting model-level generalization have been explored in terms of enhancing the training process, refinement of model architecture, or other effective optimization methods. Specifically, such strategies encompass several categories of methods: a) Learning strategy, which focuses on adequately reflecting the target-suitable knowledge or leveraging distinct representations gained from a variety of sub-tasks; b) Model framework, which exploits modifications to the network architecture or the incorporation of adaptive auxiliary components to more efficiently address the domain shift; and lastly c) Other model-based DG, which involve various optimization and adaptation techniques.
3.3.1 Learning Strategy
@@ 3.3.1.1   Meta-learning
$\phi^{*}$ $=\text{MetaLearn}(\mathcal{S}_{mtrain}),$
$\theta^{*}$ $=\text{Learn}(\mathcal{S}_{mtest};\phi^{*}),$ (13)
where $\phi^{*}$ denotes the meta-learned parameters, which are then used to learn the model parameters $\theta^{*}$ on meta-test set $\mathcal{S}_{mtest}$. The two functions, Learn($\cdot$) and MetaLearn($\cdot$), are integral components of distinct meta-learning algorithms.
@@
3.3.1.2   Self-supervised Learning
@@
3.3.1.3   Adversarial Learning
@@ 3.3.2.3   Distributed learning
3.3.3 Other Model-based DG
3.3.3.1   Geometric learning
@@
3.3.3.2   Distributionally Robust Optimization
@@
3.4 Analysis-level DG
Analysis-level domain generalization refers to the strategies and methods that aim to understand, explain, and interpret the decision-making process of machine learning models in the context of domain generalization. However, there is often a tradeoff between performance (e.g., accuracy) and interpretability, as complex and deeper models tend to be more challenging to interpret but offer better performance [127]. To address this issue, analysis-level domain generalization techniques are employed to enhance the understanding of highly generalizable models.
3.4.1 Interpretable AI
@@ 3.4.2 Causalility
4 DG Under Limited Source
TABLE IV: Comparison of Source-limted DG
@@ 4 DG Under Limited Source
In this section, we examine the challenge of domain generalization under severe restrictions in the source domain, as detailed in Table IV. These particular DG scenarios have received relatively less attention in the field of MedIA due to their extreme conditions. Our survey aims to shed light on these under-explored areas and their implications for MedIA.
4.1 Single-source Domain Generalization
[…]
4.2 Open-set Domain Generalization
[…]
4.3 Other under-explored DGs
[…]
5 Future Directions
@@
5.1 Source-limited Domain Generalization
[…]
5.2 Medical Foundation Model
@@
5.3 Benchmark Datasets
Many DG methods for MedIA rely on custom datasets that mix private and public datasets, treating each dataset as a distinct domain. The reproducibility of these custom datasets is extremely low because the processes for data splitting, preprocessing, and annotation vary widely and are hard to duplicate. Therefore, there is a need to establish standardized benchmark datasets that reflect the diversity and challenges encountered in real-world medical imaging scenarios. These benchmark datasets should cover various imaging modalities, patient populations, and imaging protocols to facilitate the fair and rigorous evaluation of domain generalization methods. Additionally, benchmark datasets for different settings of domain generalization, i.e., multi-source, single-source, cross-site, cross-sequence, cross-modality, covariate shift, and concept shift, should be developed to enable comprehensive evaluation and comparison of different techniques. Readers are refer to Appendix B in the Supplementary for a review of existing DG benchmark datasets and a summary of public dataset.
5.4 Suggestions for Domain Generalization for MedIA
[…]
There exist a number of empirical evaluations of commonly used domain generalization techniques suggesting appropriate methods for specific tasks at hand. For example, Korevaar et al[159] evaluated methods on benchmark datasets, whereas Zhang et al[160] and Galappaththige et al[147] evaluated them on some custom cross-site datasets consisting of publicly available datasets. While these benchmarks shed light on the capabilities of certain techniques in domain generalization, they do not offer a comprehensive guide for problem-specific and task-specific strategies throughout the MedIA workflow. To this end, we critically analyze different DG methods (Table LABEL:tab:critical in the Supplementary) and suggest strategies to incorporate domain generalization into the model (Fig. 4).
6 Conclusion
Domain generalization is a crucial capability for modern medical image analysis, aimed at creating machine learning models capable of handling a wide variety of data distributions arising from variations in, for example, imaging protocols, patient demographics, and equipment. This paper provides a comprehensive survey of domain generalization techniques, extending beyond the methodological hierarchy of previous surveys and considering the implications of domain generalization on the entire MedIA workflow. Our focus includes every stage of the decision-making process, from data acquisition to pre-processing, prediction, and analysis.

### v3

@@ III-B Feature-level Generalization
TABLE VI: Feature-level domain generalization methods. Methods categorized by different settings for source and target domains (see §$\S$§II-B), task, organ, and modality used in experiment.
Method Specific Ref. Source Target Task Organ Modality
Feature Alignment Feature Normalization [81] Single Sequence, Modality Segmentation Brain, Cardiac, Abdominal MRI, CT
[82] Single Site Segmentation Retinal Fundus
Dissimilarity-based [83] Multiple Site Classification Colon, Breast Histology
[84] Multiple Modality Segmentation Retinal Fundus, OCT
[85] Multiple Sequence, Site Classification, Segmentation Skin, Spinal Dermatology, MRI
Implicit Disentanglement Mutual Information [86] Multiple Site Classification Abdominal, Brain, Femur, Lips Fetal Ultrasound
[87] Multiple Site Classification Abdominal, Brain, Femur, Lips Fetal Ultrasound
[88] Single Site Segmentation Carotid Ultrasound
[89] Single, Multiple Sequence Classification Blood Cell Histology
[48] Single Modality, Sequence Segmantation Prostate, Abdominal, Cardiac MRI, CT
Contrastive [90] Multiple Site Detection Breast X-ray
[91] Multiple Site Segmentation Retinal Fundus
Variational [92] Multiple Site Classification Blood Cell Histology
[93] Multiple Site Classification Breast Histology
[94] Multiple Site Classification Breast X-ray
[95] Multiple Site Classification Breast X-ray
Explicit Disentanglement Conditional Representation Learning [96] Multiple Site Classification Retinal Fundus
[97] Multiple Site Classification Retinal Fundus
[98] Multiple Site Segmentation Retinal Fundus
Feature Reguarlization [99] Multiple Site Segmentation Cardiac MRI
[100] Multiple Site Compression, classification Brain, Skin MRI, Histology
[101] Single Site Segmentation Spinal MRI
Others Feature Augmentation [102] Single, Multiple Site Segmentation Cardiac, Prostate MRI
[103] Multiple Site Segmentation Prostate MRI
[104] Single Site Liver CT
Kernel-based [105] Single Site Classification Brain MRI
[106] Multiple Site Detection Brain EEG
@@
Refer to Table V for a summary of feature-level methods. In the following paragraphs, we explore feature-level domain generalization techniques.
III-B1 Feature Alignment
@@
III-B1b   Dissimilarity-based Alignment
@@ III-B2 Disentanglement Methods
III-B2a   Implicit Feature Disentanglement
@@
Contrastive Disentanglement aims to make representations of similar instances more alike (low contrast) and those of different instances more dissimilar (high contrast). In the context of domain generalization, this approach can be used to learn domain-invariant features by encouraging similarity between samples that share the same task-relevant characteristics, regardless of their domain, while separating samples with different characteristics. A typical contrastive learning loss function [109] is defined as:
@@
where $\operatorname{sim}(\cdot,\cdot)$ is a function for cosine similarity, $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$ are positive pairs, $k\in a(i)$ are the indexes of selected negative samples, and $\mathbf{x}_{i}$ and $\mathbf{x}_{k}$ are negative pairs. In the context of domain generalization for medical image analysis, positive pairs could be defined as images showing the same pathology (e.g., two images of malignant tumors) regardless of whether they come from different hospitals or were acquired using different imaging protocols. Negative pairs would be images showing different pathologies (e.g., an image of a malignant tumor paired with an image of healthy tissue), again regardless of their source domain. This pairing strategy encourages the model to learn features that distinguish between pathologies while being invariant to domain-specific characteristics like image acquisition settings or hospital-specific protocols.
@@
III-B2b   Explicit Feature Disentanglement
@@
III-B3 Other Representation Learning Methods
III-B3a   Feature Augmentation
@@ III-B3 Other Representation Learning Methods
III-B3b   Kernel-based Learning
Kernel-based methods are a classic and effective approach within feature-level domain generalization. These methods improve generalization by mapping the original input features into a higher dimensional space. This mapping offers several advantages for domain generalization. In the higher-dimensional space, kernel methods can potentially reveal domain-invariant structures that are not apparent in the original feature space. Kernel methods can also model non-linear relationships in the data, which is particularly useful for capturing intricate patterns in medical images that may be consistent across domains. There are various kernel-based methods for feature-level domain generalization, including Support Vector Machine (SVM) variants, Maximum Mean Discrepancy (MMD), and Transfer Component Analysis (TCA). Kernel trick enables these methods to operate in high-dimensional spaces without explicitly calculating the coordinates of the data in that space, but by simply computing the dot products between the images of all data pairs in the feature space. This makes the calculations more tractable and efficient. These kernel-based methods can benefit medical image analysis as they can handle high-dimensional data and discover complex patterns. They also offer an excellent way to incorporate domain knowledge, such as spatial relationships in images, by defining appropriate kernels.
@@
III-C Model-level Generalization
Model-level generalization focuses on enhancing the intrinsic ability of machine learning models to generalize across domains by modifying core aspects of the model itself, including the learning process, model architecture, and optimization techniques. Specifically, such strategies encompass several categories of methods: a) Learning strategy, which focuses on adequately reflecting the target-suitable knowledge or leveraging distinct representations gained from a variety of sub-tasks; b) Model framework, which exploits modifications to the network architecture or the incorporation of adaptive auxiliary components to more efficiently address the domain shift; and lastly c) Other model-based DG, which involve various optimization and adaptation techniques.
TABLE VII: Model-level domain generalization methods. Methods categorized by different settings for source and target domains, task, organ, and modality.
Method Specific Ref. Source Target Task Organ Modality
Learning Strategy Meta-learning [114] Multi Site Segmentation Spinal CT
[115] Multiple Site Segmentation Brain MRI
[116] Multiple Site Segmentation Prostate MRI
[117] Multiple Site, Sequence Segmentation Cardiac, Spinal MRI
[118] Multi Site Segmentation Retinal Fundus, OCT, FC
[119] Single, Multiple Site Classification Brain Functional MRI
Self-supervised Learning [110] Multiple Site Segmentation Retinal Fundus
[91] Multiple Site Segmentation Retinal Fundus
[120] Multiple Modality Segmentation Abdominal, Cardiac CT, MRI
[121] Multiple Site Classification Skin, Retinal, Chest, Breast Histology, Fundus, X-ray, Mammography
[122] Single Site Segmentation Retinal Fundus
Adversarial Learning [123] Single, Multiple Sequence Detection Skin, Hip MRI
[124] Multiple Site Segmentation Cardiac MRI
[125] Multiple Site Segmentation Retinal, Prostate Fundus, MRI
[126] Single Site Classification Breast Histology
Model Framework Ensemble Learning [127] Multiple Sequence, Site Segmentation Brain MRI
[128] Single Sequence Localization Surgical Scene Video Frames
[44] Single, Multiple Site Classification Colon, Breast Histology
Model Distillation [101] Single Site Segmentation Spinal MRI
[129] Multi Site Detection Breast Histology
[130] Single, Multi Site Classification Retinal Fundus
[131] Single Site Segmentation Prostate MRI
[132] Single Site Segmentation Prostate MRI
[133] Multi Site Segmentation Abdominal, Prostate, Surgical Scene CT, MRI, Video
Distributed Learning [134] Multiple Site Segmentation Retinal, Prostate Fundus, MRI
[135] Multiple Site Classification Breast Histology
[136] Multiple Site Segmentation Retinal Fundus
[137] Multiple Modality Segmentation Tumor MRI
Other Geometric Learning [138] Multiple Modal, Sequence, Site Classification, segmentation, detection Multiple organs, modalities from 55 datasets
[139] Multiple Site Segmentation Surgical Scene Video Frames
[140] Single Site Segmentation Abdominal, Cardiac, Prostate MRI, CT
Distributionally Robust Optimization [141] Multiple Site Classification Skin Dermatology
[142] Multiple Site Classification Skin Dermatology
III-C1 Learning Strategy
@@
$\phi^{*}$ $=\arg\min_{\phi}\mathcal{L}_{meta}(\phi;\mathcal{S}_{mtrain}),$
$\theta^{*}$ $=\arg\min_{\theta}\mathcal{L}_{task}(\theta;\mathcal{S}_{mtest},\phi),$ (13)
where $\phi^{*}$ denotes the meta-learned parameters optimized on the meta-training set $\mathcal{S}_{mtrain}$, which are then used to initialize the task-specific parameters $\theta^{*}$ optimized on the meta-test set $\mathcal{S}_{mtest}$. $\mathcal{L}_{meta}$ and $\mathcal{L}_{task}$ are the meta-learning and task-specific loss functions, respectively. This formulation is inspired by the Model-Agnostic Meta-Learning (MAML) algorithm [143], where the meta-objective is to find an initialization that allows for quick adaptation to new tasks.
@@
III-C1b   Self-supervised Learning
@@
An emerging and powerful strategy in addressing domain shifts is the combination of pretraining and self-supervised adaptation. This approach leverages the benefits of both large-scale pretraining and task-specific fine-tuning to enhance model generalization. A notable example is FINE (Feature-level Instance Normalization and Exchange) method proposed by Zhang et al[145], which incorporates the physical model of data generation into the adaptation process. FINE updates the weights of a pretrained network by minimizing a data fidelity loss for each test case, allowing it to better capture features specific to the target domain while maintaining physical consistency. Similarly, Zhao et al[146] introduced Synthetic Multi-Orientation Resolution Enhancement (SMORE). SMORE is a self-supervised technique for super-resolution and anti-aliasing of MRI images that does not require external training data. It works by training a network on high-resolution in-plane slices and applying it to low-resolution through-plane slices to enhance image quality.
III-C1c   Adversarial Learning
@@ III-C2 Model Framework
III-C3 Other Model-based DG
III-C3a   Geometric learning
@@ III-C3 Other Model-based DG
TABLE VIII: Analysis-level domain generalization methods. Methods categorized by different settings for source and target domains (see §$\S$§II-B), task, organ, and modality used in experiment.
Method Ref. Source Target Task Organ Modality
Interpretable AI [152] Single Site Classification, Segmentation Retinal, Cardiac OCT, CT
[153] Single Site Classification, Segmentation Knee MRI
[154] Multiple Site Classification Brain EEG
[155] Multiple Sequence Detection Gastrointestinal Endoscopy
Transferability [156] Single Sequence, Site Segmentation Cardiac, Prostate MRI, CT
[157] Single, Multiple Site Classification Breast Histology
Causality [158] Single, Multiple Sequence Classification Chest X-ray
[159] Multiple Site Classification Brain MRI
[160] Single Modality, Sequence, Site Segmentation Prostate, Abdominal, Cardiac MRI, CT
III-C3b   Distributionally Robust Optimization
@@
III-D Analysis-level Generalization
Analysis-level DG refers to techniques that focus on understanding and interpreting the behavior of domain-generalized models. These methods aim to: (1) provide insights into how models make decisions across different domains, (2) evaluate the extent of a model’s generalization, (3) identify potential biases or failure modes in generalized models, and (4) enable trust and adoption of DG models in clinical settings. Unlike data-level, feature-level, or model-level approaches that primarily aim to improve generalization performance, analysis-level methods are concerned with the post-hoc examination and interpretation of already generalized models. This is particularly crucial in MedIA, where understanding model decisions is essential for clinical validation and trust.
Analysis-level DG techniques face unique challenges, as they must provide interpretations that are consistent and meaningful across multiple domains, often with varying characteristics. These methods must balance the trade-off between model performance and interpretability, especially for complex, highly generalizable models.
III-D1 Interpretable AI
@@ III-D2 Causalility
IV DG Under Limited Source
TABLE IX: Comparison of Source-limted DG
@@ IV DG Under Limited Source
In this section, we examine the challenge of domain generalization under severe restrictions in the source domain, as detailed in Table IX. These particular DG scenarios have received relatively less attention in the field of MedIA due to their extreme conditions. Our review aims to shed light on these under-explored areas and their implications for MedIA.
IV-A Single-source Domain Generalization
Single-source domain generalization (SSDG) assumes that there is only one source domain to learn from. Due to the lack of diversity in the training data, most SSDG propose augmentation-based solutions, both in data and feature space, to simulate a broader range of domain variability. For example, augmentation-based SSDG inlcude color normalization [30] (§$\S$§III-A1a), frequency-based [35] (§$\S$§III-A1b), randomization [41, 42] (§$\S$§III-A2a), generative models [46] (§$\S$§III-A2c), feature augmentation [104] (§$\S$§III-B3a), contrastive learning [122] (§$\S$§III-C1b), adversarial learning [48, 126] (§$\S$§III-C1c). Recent bleeding-edge SSDG include transferring knowledge from large-scale pretrained models, i.e. variations of model distillation [130, 131, 132] (§$\S$§III-C2b).
IV-B Open-set Domain Generalization
TABLE X: Open-set domain generalization methods. Methods categorized by different settings for source and target domains (see §$\S$§II-B), task, organ, and modality used in experiment.
Method Ref. Source Target Task Organ Modality
Open-set DG [166] Single Site Classification Chest X-ray
[167] Multiple Site Classification Chest X-ray
[168] Single Site Classification Chest X-ray
[169] Single Site Classification Chest X-ray
[170] Single Site Classification Chest X-ray
[171] Multiple Site Classification Breast Histology
Open-set domain generalization (OSDG) refers to DG techniques that specialize in capturing and correcting the concept shift in addition to the covariate shift (§$\S$§II-B3). Yang et al[166] proposed a simple feature-based semantics score function to consider both detecting label shift and being tolerant to covariate shift as in-distribution. Mahajan et al[167] investigated the theoretical relationship of whether better OOD generalization leads to better privacy for ML models in practice and showed that capturing stable features from models represents superior open-set generalization with robustness. Zheng et al[168] proposed Open-Set Single-Domain Generalization for Multiple Cross-Matching (MCM) for the open-set lung cancer diagnosis. This work delves into an open-set single-source DG problem where the source domain only contains data with unique class names, while the target domain contains multiple unseen class names. Puli et al[169, 170] and Gao et al[171] deal with the spurious correlation or variations underlying several confounding variables in terms of causal perspective to circumvent the open-set problem.
IV-C Other under-explored DGs
In this section, we introduce some under-explored DGs where only preliminary research has been done in the field of MedIA. Unsupervised domain generalization refers to methods that enable a model to learn useful, domain-invariant features from unlabeled source data such that it can perform well on unseen domains. Source-free domain generalization places an extreme constraint on privacy-preserving models (§$\S$§III-C2c) where it assumes source data is inaccessible, but only the pretrained source model is available. Zero-shot domain generalization is another extreme case of DG where only the auxiliary information (e.g., meta-data) from the source domain is available. A possible solution [130, 131, 132, 133] (§$\S$§III-C2b) to these extreme cases of DGs involves a pretrained foundation model combined with zero-shot prompting techniques, though these topics fall beyond the scope of this survey due to the lack of literature in this topic.
V Future Directions
§$\S$§III-A1b Frequency-based Start §$\S$§II-C Related ML tasks Sufficient data? Raw measurement available? Access to $\mathcal{S}_{target}$? §$\S$§III-A2 Data Augmentation Preprocessing required? Specific task or problem at hand? Analysis or interpretability required? §$\S$§III-A3 Problem-specific §$\S$§III-A1a Image Processing §$\S$§III-D Analysis-level DG Able to use special algorithm or architecture? §$\S$§III-C Model-level DG §$\S$§III-B1 Feature Alignment Invariant v.s. Disentangled Features §$\S$§III-B2 Feature Disentanglement YNYNYNYYNNYNYNInv.Dis.
TABLE XI: Critical Analysis of Domain Generalization Methods
Level Method Specific Strengths Limitations
Data Image Process. Intensity Normalization Uniformity across images; improved ML performance May reduce contrast; relies on similarity assumptions
Histogram Matching Adapts to reference histograms improving consistency Depends on reference choice; less suitable for multi-domain
Color Normalization Standardizes color for consistency and feature recognition Could alter diagnostic features; mostly for stain images
Surrogate Frequency-based Separates amplitude/phase for style/content manipulation Sensitive to noise; limited by domain applicability
Raw Signals Manages early-stage data manipulation; captures latent info Complexity/accessibility issues; limited applicability
Dictionary Learning Captures sparse representation and structural similarities Computationally intense; may not handle unique features
Augment Randomization Increases data diversity and simplicity Limited augmentation scope; may create unrealistic images
Adversarial Improves robustness through targeted examples Computationally intense; risks catastrophic forgetting
Generative Generates diverse and realistic data Complex to train; prone to modal-collapse
Problem-specific Cross-modal Enhances unpaired data utilization Complex and relies on synthetic data quality
Stain Normalization Enhances histopathological analysis Limited applicability; risks of over-normalization
Feature Feature Align. Feature Normalization Standardizes statistical distributions efficiently Depends on choice of domain-specific statistics
Dissimilarity-based Mitigates domain shift through direct measures Computationally complex and challenges in metric selection
Implicit Disent. Information Theoretic Enhances interpretability and adapts to complex structures Difficult to estimate MI; depends on estimation quality
Contrastive Improves sample efficiency; robustness to domain shifts Depends on pair quality; risks representational collapse
Variational Models uncertainty; direct latent space regularization Balance between fidelity and disentanglement
Explicit Disent. Conditional Representation Enhances contextual adaptation; targeted feature learning Complexity leading to potential overfitting on condition
Feature Regularization Robustness to variations; prevents over-reliance on features Sensitive to choice of hyperparameters
Others Feature Augmentation Incorporates domain knowledge within latent space Complex and can overfit to augmented features
Kernel-based Learning Incorporates domain knowledge efficiently Selection of kernel is critical; scalability issues
Model Learning Strategy Meta-learning Enables rapid adaptation and efficient learning Faces complex optimization; overfitting to meta-tasks
Self-supervised Learning Utilizes unlabeled data for enhanced features Dependent on pretext task design
Adversarial Learning Learns domain-invariant features and enhances robustness Vulnerable to adversarial attacks; computationally complex
Model Framework Ensemble Learning Increases robustness and leverages model diversity High computational cost and implementation complexity
Distillation Highly efficient; preserves ID performance Dependent on choice of teacher; complex training process
Distributed Learning Preserves privacy and improves scalability Faces communication overhead and heterogeneity issues
Others Geometric Handles non-Euclidean data well High complexity; limited applicability
DRO Optimizes across worst-case scenarios for robustness Complex in defining uncertainty sets
Analysis Interpretable AI Enhances trust; model debugging Interpretability-complexity tradeoff
Causality Focuses on invariant features for robust generalization Complex model development and limited scalability
@@ V Future Directions
V-A Source-limited Domain Generalization
Unlike the common focus on scenarios involving multiple sources, there is a significant gap in research for methods tailored to domain generalization under limited source (§$\S$§IV). For example, single-source domain generalization requires the model to generalize well to unseen domains even when only a single source domain is available for training. This scenario often arises in medical imaging, where data collection can be resource-intensive, and privacy concerns may limit access to multiple sources. Future research should explore novel methods that effectively address the challenges of source-limited domain generalization, such as robustness to concept shift in the presence of covariate shift.
V-B Medical Foundation Model
@@
V-C Benchmark Datasets
Many DG methods for MedIA rely on custom datasets that mix private and public datasets, treating each dataset as a distinct domain. The reproducibility of these custom datasets is extremely low because the processes for data splitting, preprocessing, and annotation vary widely and are hard to duplicate. Therefore, there is a need to establish standardized benchmark datasets that reflect the diversity and challenges encountered in real-world medical imaging scenarios. These benchmark datasets should cover various imaging modalities, patient populations, and imaging protocols to facilitate the fair and rigorous evaluation of domain generalization methods. Additionally, benchmark datasets for different settings of domain generalization, i.e., multi-source, single-source, cross-site, cross-sequence, cross-modality, covariate shift, and concept shift, should be developed to enable comprehensive evaluation and comparison of different techniques. Readers are referred to the Supplementary and the interactive website (https://milab.korea.ac.kr/dg-dataset) for a review of existing DG benchmark datasets and a summary of public datasets.
V-D Suggestions for Domain Generalization for MedIA
There exist a number of empirical evaluations of commonly used domain generalization techniques suggesting appropriate methods for specific tasks at hand. For example, Korevaar et al[174] evaluated methods on benchmark datasets, whereas Zhang et al[175] and Galappaththige et al[130] evaluated them on some custom cross-site datasets consisting of publicly available datasets. While these benchmarks shed light on the capabilities of certain techniques in domain generalization, they do not offer a comprehensive guide for problem-specific and task-specific strategies throughout the MedIA workflow. To this end, we critically analyze different DG methods (Table XI) and suggest strategies to incorporate domain generalization into the model (Fig. 5).
VI Conclusion
Domain generalization is a crucial capability for modern medical image analysis, aimed at creating machine learning models capable of handling a wide variety of data distributions arising from variations in, for example, imaging protocols, patient demographics, and equipment. This paper provides a comprehensive review of domain generalization techniques, extending beyond the methodological hierarchy of previous surveys and considering the implications of domain generalization on the entire MedIA workflow. Our focus includes every stage of the decision-making process, from data acquisition to pre-processing, prediction, and analysis.